OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

what is slope?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

x+4y=6 4y=-x+6 y=-1/4x+3/2 now what is slope parallel lines have same slope

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

no -1/4

OpenStudy (anonymous):

oh why?

OpenStudy (anonymous):

y=mx +c m is slope i.e co efficint of x

OpenStudy (anonymous):

oh ok so i write it out how u gave it to me der?

OpenStudy (anonymous):

as slope of parallel lines are same so slope =-1/4 passing through (-8, 5). equ ?

OpenStudy (anonymous):

y-5=-1/4x+2